# Ctrl-Adapter: An Efficient and Versatile Framework for Adapting Diverse Controls to Any Diffusion Model

### v1

@@ Abstract
ControlNets are widely used for adding spatial control in image generation with different conditions, such as depth maps, canny edges, and human poses. However, there are several challenges when leveraging the pretrained image ControlNets for controlled video generation. First, pretrained ControlNet cannot be directly plugged into new backbone models due to the mismatch of feature spaces, and the cost of training ControlNets for new backbones is a big burden for many users. Second, ControlNet features for different frames might not effectively handle the temporal consistency of objects. To address these challenges, we introduce Ctrl-Adapter, an efficient and versatile framework that adds diverse controls to any image/video diffusion models, by adapting pretrained ControlNets (and improving temporal alignment for videos). Ctrl-Adapter provides strong and diverse capabilities including image control, video control, video control with sparse frames, multi-condition control, compatibility with different backbone models, adaptation to unseen control conditions, and video editing. In the Ctrl-Adapter framework, we train adapter layers that fuse pretrained ControlNet features to different image/video diffusion models, while keeping the parameters of the ControlNets and the diffusion models frozen. Ctrl-Adapter consists of temporal as well as spatial modules so that it can effectively handle the temporal consistency of videos. Additionally, for robust adaptation to different backbone models and sparse control, we propose latent skipping and inverse timestep sampling. Moreover, Ctrl-Adapter enables control from multiple conditions by simply taking the (weighted) average of ControlNet outputs. From our experiments with diverse image and video diffusion backbones (SDXL, Hotshot-XL, I2VGen-XL, and SVD), Ctrl-Adapter matches ControlNet on the COCO dataset for image control, and even outperforms all baselines for video control (achieving the state-of-the-art accuracy on the DAVIS 2017 dataset) with significantly lower computational costs (Ctrl-Adapter outperforms baselines in less than 10 GPU hours). Lastly, we provide comprehensive ablations for our design choices and qualitative examples.
Figure 1: We propose Ctrl-Adapter, an efficient and versatile framework for adding diverse spatial controls to any image or video diffusion model. Ctrl-Adapter supports a variety of useful applications, including video control, video control with multiple conditions, video control with sparse frame conditions, image control, zero-shot transfer to unseen conditions, and video editing.
@@ 1 Introduction
Recent diffusion models have achieved significant progress in generating high-fidelity images [52, 44, 56, 47] and videos [3, 15, 6, 35, 37] from text descriptions. As it is often hard to describe every image/video detail only with text, there have been many works to control diffusion models in a more fine-grained manner by providing additional condition inputs such as bounding boxes [33, 71], reference object images [54, 14, 31], and segmentation maps [13, 2, 74]. Among them, Zhang et al[74] have released a variety of ControlNet checkpoints based on Stable Diffusion [52] v1.5 (SDv1.5), and the user community has also shared many ControlNets trained with different input conditions. This has made ControlNet [74] one of the most popular spatial control methods for image generation.
@@
However, there are challenges when using the existing pretrained image ControlNets for controllable video generation. First, pretrained ControlNet cannot be directly plugged into new backbone models, and the cost for training ControlNets for new backbone models is a big burden for many users due to high computational costs. For example, training a ControlNet for SDv1.5 takes 500-600 A100 GPU hours.111 https://huggingface.co/lllyasviel/sd-controlnet-depth and https://huggingface.co/lllyasviel/sd-controlnet-canny Second, ControlNet is designed for controllable image generation; hence applying these pretrained ControlNets with image backbones directly to each video frame independently does not take the temporal consistency across video frames into account. This leads to an interesting question: Can we design an efficient and versatile framework that can reuse and adapt any existing pretrained ControlNet to accurately guide new image/video diffusion models and support diverse controls?
@@
To address this challenge, we design Ctrl-Adapter, a novel, flexible framework that enables the efficient reuse of pretrained ControlNets for diverse controls with any new image/video diffusion models, by adapting pretrained ControlNets (and improving temporal alignment for videos). We illustrate the overall capabilities of Ctrl-Adapter framework in Fig. 1. As shown in Fig. 2 left, Ctrl-Adapter trains adapter layers [26, 72] to map the features of a pretrained image ControlNet to a target image/video diffusion model, while keeping the parameters of the ControlNet and the backbone diffusion model frozen. As shown in Fig. 2 right, each Ctrl-Adapter consists of four modules: spatial convolution, temporal convolution, spatial attention, and temporal attention. The temporal convolution/attention modules effectively fuse the ControlNet features into image/video diffusion models for better temporal consistency. Additionally, to ensure robust adaptation of ControlNets to backbone models of different noise scales and sparse frame control conditions, we propose skipping the visual latent variable from the ControlNet inputs. We also introduce inverse timestep sampling to effectively adapt ControlNets to new backbones equipped with continuous diffusion timestep samplers. For more accurate control beyond a single condition, Ctrl-Adapter’s reuse of pretrained ControlNets also allows easy combination of multiple conditions (e.g., using both depth map and canny edge as conditions) by simply taking the (weighted) average of different ControlNet outputs.
@@
As shown in Table 1, Ctrl-Adapter allows many useful capabilities including image control, video control, video control with sparse frames, multi-condition control, and compatibility with different backbone models, while previous methods only support a small subset of them (see details in Sec. 2). We demonstrate the effectiveness of Ctrl-Adapter through extensive experiments and analyses. It exhibits strong performance when adapting ControlNets (pretrained with SDv1.5) to various video and image diffusion backbones, including text-to-video generation (Hotshot-XL [41]), image-to-video generation (I2VGen-XL [75] and Stable Video Diffusion [3]), and text-to-image generation (SDXL [44]).
Table 1: Overview of the capabilities supported by recent methods for controlled image/video generation. Ctrl-Adapter supports five useful features, while previous methods support only some of them.
@@
In our experiments, we first demonstrate that Ctrl-Adapter matches the performance of a pretrained image ControlNet on COCO dataset [36] and outperforms previous methods in controllable video generation (achieving state-of-the-art performance on the DAVIS 2017 dataset [45]) with significantly lower computational costs (Ctrl-Adapter outperforms baselines in less than 10 GPU hours). Next, we demonstrate that Ctrl-Adapter enables more accurate video generation with multiple conditions compared to a single condition. Averaging ControlNet features with learnable linear weights enhances performance beyond that achieved by averaging with equal weights. In addition, we show that skipping the visual latent variable from ControlNet inputs allows video control only with a few frames of (i.e., sparse) conditions, eliminating the need for dense conditions across all frames. We also highlight zero-shot adaption – Ctrl-Adapter trained with one condition (e.g., depth maps) can easily adapt to another ControlNet trained with a different condition (e.g., segmentation maps). Lastly, we provide comprehensive ablations for Ctrl-Adapter design choices and qualitative examples.
@@
2 Related Works
Text-to-video and image-to-video generation models.
Adding control to image/video diffusion models.
While recent image/video diffusion models demonstrate impressive performance in generating highly realistic images/videos from text descriptions, it is hard to describe every detail of images/videos only with text or first frame image. Instead, there have been many works using different types of additional inputs to control the image/video diffusion models, such as bounding boxes [33, 71], reference object image [54, 14, 31], segmentation map [13, 2, 74], sketch [74], etc., and combinations of multiple conditions [29, 46, 78, 66]. As finetuning all the parameters of such image/video diffusion models is computationally expensive, several methods, such as ControlNet [74], have been proposed to add conditional control capability via parameter-efficient training [74, 55, 40]. X-Adapter [48] learns an adapter module to reuse ControlNets pretrained with a smaller image diffusion model (e.g., SDv1.5) for a bigger image diffusion model (e.g., SDXL). While they focus solely on learning an adapter for image control, Ctrl-Adapter features architectural designs (e.g., temporal convolution/attention layers) for video generation as well. In addition, X-Adapter needs the smaller image diffusion model (SDv1.5) during training and inference, whereas Ctrl-Adapter doesn’t need the smaller diffusion model at all (for image/video generation), hence being more memory and computationally efficient (see Appendix C for details). SparseCtrl [17] guides a video diffusion model with conditional inputs of few frames (instead of full frames), to alleviate the cost of collecting video conditions. Since SparseCtrl involves augmenting ControlNet with an additional channel for frame masks, it requires training a new variant of ControlNet from scratch. In contrast, we leverage existing image ControlNets more efficiently by propagating information through temporal layers in adapters and enabling sparse frame control via skipping the latents from ControlNet inputs (see Sec. 3.2 for details). Furthermore, compared with previous works that are specially designed for specific condition controls on a single modality (image [74, 46] or video [27, 76]), our work presents a unified and versatile framework that supports diverse controls, including image control, video control, sparse frame control, and multi-source control, with significantly lower computational costs by reusing pretrained ControlNets (e.g., Ctrl-Adapter outperforms baselines in less than 10 GPU hours). Table 1 summarizes the comparison of Ctrl-Adapter with related works.
@@ 3 Method
In the section, we first explain the background of latent diffusion models and ControlNets (Sec. 3.1). Next, we introduce Ctrl-Adapter, a novel framework that enables the efficient reuse of pretrained ControlNets for diverse controls with any new image/video diffusion models (Sec. 3.2). Our Ctrl-Adapter matches or outperforms existing image/video control methods in both full-frame and sparse-frame condition settings, achieving state-of-the-art video control performance, with significantly lower training costs (e.g., Ctrl-Adapter trained for 22 hours can match SDXL ControlNet trained for 700 hours; see Fig. 13). Lastly, we explain image/video generation with multiple conditions (e.g., depth/human pose) can be achieved by combining multiple ControlNets with different input conditions via a simple (weighted) averaging (Sec. 3.3).
@@ 3.2 Ctrl-Adapter
We introduce Ctrl-Adapter, a novel framework that enables the efficient reuse of existing image ControlNets (of SDv1.5) for spatial control with new diffusion models. We mainly describe our method details in the video generation settings, since Ctrl-Adapter can be flexibly adapted to image diffusion models by regarding images as single-frame videos.
Efficient Adaptation of Pretrained ControlNets.
@@
As shown in Fig. 2 (left), we train an adapter module (colored orange) to map the middle/output blocks of a pretrained ControlNet (colored blue) to the corresponding middle/output blocks of the target video diffusion model (colored green). If the target backbone does not have the same number of output blocks Ctrl-Adapter maps the ControlNet features to the output block that handles the closest height and width of the latents. We keep all parameters in both the ControlNet and the target video diffusion model frozen. Therefore, training a Ctrl-Adapter can be significantly more efficient than training a new video ControlNet.
Ctrl-Adapter architecture.
@@
As shown in Fig. 2 (right), each block of Ctrl-Adapter consists of four modules: spatial convolution, temporal convolution, spatial attention, and temporal attention. We set the values for $N_{1:4}$ and $N$ as 1 by default. All four modules aim to align the image ControlNet features with those of target video diffusion models. And the temporal convolution and attention modules model effectively fuse the in ControlNet features for better temporal consistency. Moreover, the spatial/temporal convolution modules incorporate the current denoising timestep $t$ and spatial/temporal convolution modules incorporate the conditions (i.e., text prompt/initial frame) $\bm{c}_{\text{text/img}}$. This design allows the Ctrl-Adapter to dynamically adjust its features according to different denoising stages and the objects generated. When adapting to image diffusion models, Ctrl-Adapter blocks only consist of spatial convolution/attention modules (without temporal convolution/attention modules). See Section 6.1 for detailed ablation studies on the design choices of Ctrl-Adapter, and see Appendix B for the implementation details of the four modules.
Figure 3: Left: latent $\bm{z}_{t}$ is given to ControlNets (default Ctrl-Adapter). Right: latent $\bm{z}_{t}$ is not given to ControlNet.
Skipping the latent from ControlNet inputs: robust adaption to different noise scales & sparse frame conditions.
@@
Although the original ControlNets take the latent $\bm{z}_{t}$ as part of their inputs, we find that skipping $\bm{z}_{t}$ from ControlNet inputs is effective for Ctrl-Adapter in certain settings, as illustrated in Fig. 3. (1) Different noise scales: while SDv1.5 samples noise $\bm{\epsilon}$ from $N(\bm{0},\bm{I})$, some recent diffusion models [25, 10, 3] sample noise $\bm{\epsilon}$ of much bigger scale (e.g. SVD [3] sample noise from $\sigma*N(\bm{0},\bm{I})$, where $\sigma\sim\text{LogNormal}(0.7,1.6)$; $\sigma\in[0,+\infty]$ and $\mathbb{E}[\sigma]=7.24$). We find that adding larger-scale $\bm{z}_{t}$ from the new backbone models to image conditions $\bm{c}_{f}$ dilutes the $\bm{c}_{f}$ and makes the ControlNet outputs uninformative, whereas skipping $\bm{z}_{t}$ enables the adaptation of such new backbone models. (2) Sparse frame conditions: when the image conditions are provided only for the subset of video frames (i.e., $\bm{c}_{f}=\emptyset$ for most frames $f$), ControlNet could rely on the information from $\bm{z}_{t}$ and ignore $\bm{c}_{f}$ during training. Skipping $\bm{z}_{t}$ from ControlNet inputs also helps the Ctrl-Adapter to more effectively handle such sparse frame conditions (see Table 5).
Inverse timestep sampling: robust adaptation to continuous diffusion timestep samplers.
@@
While SDv1.5 samples discrete timesteps $t$ uniformly from $\{0,1,...1000\}$, some recent diffusion models [10, 38, 51] sample timesteps from continuous distributions, e.g., SVD [3] first samples noise scale $\sigma_{\text{SVD}}$ from a log-normal distribution $\sigma_{\text{SVD}}\sim\text{LogNormal}(0.7,1.6)$, and then sets the timestep $t_{\text{SVD}}$ as a function of $\sigma_{\text{SVD}}$: $t_{\text{SVD}}=0.25\log\sigma_{\text{SVD}}$. This gap between discrete and continuous distributions means that we cannot assign the same timestep $t$ to both the video diffusion model and the ControlNet. Therefore, we propose inverse timestep sampling, an algorithm that creates a mapping between the continuous and discrete time distributions to match different timestep samplers (see Algorithm 1 for PyTorch [1] code). During each training step, the procedure for this algorithm can be summarized as follows: (1) we sample a variable $u$ from $\text{Uniform}[0,1]$; (2) we sample noise scale $\sigma_{\text{cont.}}$ via inverse transform sampling; i.e., we derive the inverse cumulative density function of $\sigma_{\text{cont.}}$ and sample $\sigma_{\text{cont.}}$ by sampling $u$: $\sigma_{\text{cont.}}=F_{\text{cont.}}^{-1}(u)$; (3) given a preconditioning function $g_{\text{cont.}}$ that maps noise scale to timestep (typically associated with the continuous-time noise sampler), we can compute $t_{\text{cont.}}=g_{\text{cont.}}(\sigma_{\text{cont.}})$; (4) we set the timesteps and noise scales for both ControlNet and our Ctrl-Adapter as $t_{\text{CNet}}=1000u$ and $\sigma_{\text{CNet}}=u$ respectively, where 1000 represents the denoising timestep range over which the ControlNet is trained. During inference, we uniformly sample $k$ equidistant values for $u$ within $[0,1]$ and derive corresponding $t_{\text{cont./CNet}}$ and $\sigma_{\text{cont./CNet}}$ as inputs for denoising steps.
Figure 4: Multi-condition video generation by combining multiple ControlNets.

### v2

@@ Abstract
ControlNets are widely used for adding spatial control to text-to-image diffusion models with different conditions, such as depth maps, scribbles/sketches, and human poses. However, when it comes to controllable video generation, ControlNets cannot be directly integrated into new backbones due to feature space mismatches, and training ControlNets for new backbones can be a significant burden for many users. Furthermore, applying ControlNets independently to different frames cannot effectively maintain object temporal consistency. To address these challenges, we introduce Ctrl-Adapter, an efficient and versatile framework that adds diverse controls to any image/video diffusion model through the adaptation of pretrained ControlNets. Ctrl-Adapter offers strong and diverse capabilities, including image and video control, sparse-frame video control, fine-grained patch-level multi-condition control (via an MoE router), zero-shot adaptation to unseen conditions, and supports a variety of downstream tasks beyond spatial control, including video editing, video style transfer, and text-guided motion control. With six diverse U-Net/DiT-based image/video diffusion models (SDXL, PixArt-$\alpha$, I2VGen-XL, SVD, Latte, Hotshot-XL), Ctrl-Adapter matches the performance of pretrained ControlNets on COCO and achieves the state-of-the-art on DAVIS 2017 with significantly lower computation ($<$ 10 GPU hours).
Figure 1: We propose Ctrl-Adapter, an efficient and versatile framework for adding diverse controls to any diffusion model. Ctrl-Adapter supports a variety of useful applications.
@@ 1 Introduction
Recent diffusion models have achieved significant progress in generating high-fidelity images [61, 52, 65, 55] and videos [3, 18, 6, 41, 43] from text descriptions. As it is often hard to describe every image/video detail only with text, there have been many works to control diffusion models in a more fine-grained manner by providing additional condition inputs such as bounding boxes [39, 83], reference object images [63, 17, 36], and segmentation maps [16, 2, 86]. Among them, Zhang et al[86] have released a variety of ControlNet checkpoints based on Stable Diffusion [61] v1.5 (SDv1.5), and the user community has shared many ControlNets trained with different input conditions. Until now, ControlNet has become one of the most popular methods for controllable image generation.
@@
However, there are challenges when using the existing pretrained image ControlNets for controllable video generation. First, pretrained ControlNeta cannot be directly plugged into new backbone models, and the cost for training ControlNets for new backbone models is a big burden for many users due to high computational costs. For example, training a ControlNet for SDv1.5 takes 500-600 A100 GPU hours [87, 88]. Second, ControlNet was originally designed for controllable image generation; hence, applying pretrained image ControlNets directly to each video frame independently does not take the temporal consistency across frames into account.
@@
To address this challenge, we design Ctrl-Adapter, a novel, flexible framework that enables the efficient reuse of pretrained ControlNets for diverse controls with any new image/video diffusion models, by adapting pretrained ControlNets (and improving temporal alignment for videos). We illustrate the overall capabilities of Ctrl-Adapter framework in Fig. 1. As shown in Fig. 3 left, Ctrl-Adapter trains adapter layers [30, 84] to map the features of a pretrained image ControlNet to a target image/video diffusion model, while keeping the parameters of the ControlNet and the backbone diffusion model frozen. As shown in Fig. 3 right, each Ctrl-Adapter consists of four modules: spatial convolution, temporal convolution, spatial attention, and temporal attention. The temporal convolution/attention modules effectively fuse the ControlNet features into image/video diffusion models for better temporal consistency. Additionally, to ensure robust adaptation of ControlNets to backbone models of different noise scales and sparse frame control conditions, we propose skipping the visual latent variable from the ControlNet inputs. We also introduce inverse timestep sampling to effectively adapt ControlNets to new backbones equipped with continuous diffusion timestep samplers. For more accurate control beyond a single condition, we designed a novel and powerful Mixture-of-Experts (MoE) router, which allows fine-grained, patch-level composition of spatial feature maps from multiple control conditions via Ctrl-Adapters (see Sec. 3.3).
@@
As shown in Table 1, Ctrl-Adapter allows many useful capabilities, including image control, video control, video control with sparse frames, multi-condition control, and compatibility with different backbone models, while previous methods only support a small subset of them (see details in Sec. 2). We demonstrate the effectiveness of Ctrl-Adapter through extensive experiments and analyses. It exhibits strong performance when adapting ControlNets (pretrained with SDv1.5) to various video and image diffusion backbones, including image-to-video generation – I2VGen-XL [89] and Stable Video Diffusion (SVD) [3], text-to-video generation – Latte [46] and Hotshot-XL [49], and text-to-image generation – SDXL [52] and PixArt-$\alpha$ [8]. The ability of Ctrl-Adapter to seamlessly adapt to DiT-based models such as Latte and PixArt-$\alpha$, which are structurally different from U-Net based ControlNets, demonstrates the flexibility of our framework design.
In Sec. 5.1 and Sec. 5.2, we first show that Ctrl-Adapter matches the performance of a pretrained image ControlNet on COCO dataset [42] and outperforms previous methods in controllable video generation (achieving state-of-the-art performance on the DAVIS 2017 dataset [53]) with significantly lower training costs (less than 10 GPU hours, see Fig. 2). Next, we demonstrate that Ctrl-Adapter enables more accurate video generation with multiple conditions compared to a single condition. Our fine-grained patch-level MoE router consistently outperforms both the equal weights baseline and the global weights MoE router (Sec. 5.3). In addition, we show that skipping the visual latent variable from ControlNet inputs allows video control only with a few frames of (i.e., sparse) conditions, eliminating the need for dense conditions across all frames (Sec. 5.4). We also highlight zero-shot adaption – Ctrl-Adapter trained with one condition can easily adapt to another ControlNet trained with a different condition (Sec. 5.5). Moreover, our Ctrl-Adapter can be flexibly applied to a variety of downstream tasks beyond spatial control, including video editing, video style transfer, and text-guided object motion control (Sec. 6). Lastly, we provide comprehensive ablations for Ctrl-Adapter design choices and qualitative examples (Appendix E, Appendix F, and Appendix G).
Table 1: Overview of the capabilities supported by controllable image/video generation methods.
@@
Figure 2: Training speed of Ctrl-Adapter for video (left) and image (right) control with depth maps, measured on A100 80GB GPUs. For both video and image controls, Ctrl-Adapter trained for 10 GPU hours outperforms strong baselines, including SDXL, which is trained for 700 GPU hours.
2 Related Works: Adding Control to Diffusion Models
There have been many works using different types of additional inputs to control the image/video diffusion models, such as bounding boxes [39, 83], reference object image [63, 17, 36], segmentation map [16, 2, 86], sketch [86], etc., and combinations of multiple conditions [33, 54, 92, 77]. As finetuning all the parameters of such image/video diffusion models is computationally expensive, several methods, such as ControlNet [86], have been proposed to add conditional control capability via parameter-efficient training [86, 64, 48].
X-Adapter [56] learns an adapter module to reuse ControlNets pretrained with a smaller image diffusion model (e.g., SDv1.5) for a bigger image diffusion model (e.g., SDXL). While they focus solely on learning an adapter for image control, Ctrl-Adapter features architectural designs (e.g., temporal convolution/attention layers) for video generation as well. In addition, X-Adapter needs to be used with the source image diffusion model (SDv1.5) during both training and inference, whereas Ctrl-Adapter does not require the smaller diffusion model for image or video generation, making it more memory and computationally efficient (see Sec. B.3 for details).
SparseCtrl [21] guides a video diffusion model with conditional inputs of few frames (instead of full frames), to alleviate the cost of collecting video conditions. Since SparseCtrl involves augmenting ControlNet with an additional channel for frame masks, it requires training a new variant of ControlNet from scratch. In contrast, we leverage existing image ControlNets more efficiently by propagating information through temporal layers in adapters and enabling sparse frame control via skipping the latents from ControlNet inputs (see Sec. 3.2 for details).
Furthermore, compared with previous works that are specially designed for specific condition controls on a single modality (image [86, 54] or video [31, 90]), our work presents a unified and versatile framework that supports diverse controls, including image control, video control, sparse frame control, with significantly lower computational costs by reusing pretrained ControlNets (outperforms strong baselines in less than 10 GPU hours, see Fig. 2). To the best of our knowledge, we are also the first work that extends multi-condition video control into fine-grained patch-level composition. Table 1 compares Ctrl-Adapter with other relevant methods. See Sec. A.1 for extended related works.
@@ 3.2 Ctrl-Adapter
We introduce Ctrl-Adapter, a novel framework that enables the efficient reuse of existing image ControlNets (SDv1.5) for spatial control with new diffusion models. We mainly describe our method details in the video generation settings, since Ctrl-Adapter can be flexibly adapted to image diffusion models by regarding images as single-frame videos.
@@
Efficient adaptation of pretrained ControlNets. As shown in Fig. 3 (left), we train an adapter module (colored orange) to map the middle/output blocks of a pretrained ControlNet (colored blue) to the corresponding middle/output blocks of the target video diffusion model (colored green). If the target backbone does not have the same number of output blocks Ctrl-Adapter maps the ControlNet features to the output block that handles the closest height and width of the latents. We keep all parameters in both the ControlNet and the target video diffusion model frozen. Therefore, training a Ctrl-Adapter can be significantly more efficient than training a new video ControlNet.
@@
Ctrl-Adapter architecture. As shown in Fig. 3 (right), each block of Ctrl-Adapter consists of four modules: spatial convolution, temporal convolution, spatial attention, and temporal attention. We set the values for $N_{1},...,N_{4}$ and $N$ as 1 by default. The temporal convolution and attention modules effectively fuse the ControlNet features to the video backbone models for better temporal consistency. Moreover, the spatial/temporal convolution modules incorporate the current denoising timestep $t$ and spatial/temporal attention modules incorporate the conditions (i.e., text prompt/initial frame) $\bm{c}_{\text{text/img}}$. This design allows Ctrl-Adapter to dynamically adjust its features according to different denoising stages and the objects generated. In addition, we skip the temporal convolution/attention modules when adapting to image diffusion models. See Sec. B.1 for architecture details of the four modules, and Appendix E for detailed ablation studies on the design choices of Ctrl-Adapter.
Adaptation to DiT-based image/video backbones. Our Ctrl-Adapter can also adapt U-Net based ControlNets to DiT-based image/video generation backbones. One important observation we made is that the spatial features encoded in the U-Net of ControlNets and the DiT blocks are structurally different (see Fig. 22). Specifically, the representation from U-Net blocks exhibits coarse-to-fine, hierarchical patterns (e.g., earlier blocks output smaller size feature maps and control high-level information such as object presence, while later blocks output larger feature maps and control lower-level details like textures), while all DiT blocks handle the feature maps of same sizes. This indicates that mapping all middle/output blocks of ControlNet to DiT blocks might not be the optimal solution. Therefore, we choose to map the feature maps of the largest size in ControlNet (i.e., block A) to the DiT blocks via Ctrl-Adapters, which are followed by zero-convolutions for channel dimension matching. To improve computational efficiency for DiT-based video generation models (i.e., Latte [46]), we only insert Ctrl-Adapters into every other DiT block (i.e., blocks 2, 4, 6…, 28, see (a) in Fig. 17). See Sec. E.2 for more discussion on Ctrl-Adapter designs for DiT.
Figure 4: Left (default): latent $\bm{z}_{t}$ is given to ControlNet. Right: latent $\bm{z}_{t}$ not given to ControlNet.
@@
Skipping the latent from ControlNet inputs: robust adaption to different noise scales & sparse frame conditions. Although the original ControlNets take the latent $\bm{z}_{t}$ as part of their inputs, we find that skipping $\bm{z}_{t}$ from ControlNet inputs is effective for Ctrl-Adapter in certain settings, as illustrated in Fig. 4. (1) Different noise scales: while SDv1.5 samples noise $\bm{\epsilon}$ from $N(\bm{0},\bm{I})$, some recent diffusion models [29, 12, 3] sample noise $\bm{\epsilon}$ of much bigger scale (e.g. SVD [3] sample noise from $\sigma*N(\bm{0},\bm{I})$, where $\sigma\sim\text{LogNormal}(0.7,1.6)$; $\sigma\in[0,+\infty]$ and $\mathbb{E}[\sigma]=7.24$). We find that adding larger-scale $\bm{z}_{t}$ from the new backbone models to image conditions $\bm{c}_{f}$ dilutes the $\bm{c}_{f}$ and makes the ControlNet outputs less informative, whereas skipping $\bm{z}_{t}$ enables the adaptation of such new backbone models. (2) Sparse frame conditions: when the image conditions are provided only for the subset of video frames (i.e., $\bm{c}_{f}=\emptyset$ for most frames $f$), ControlNet could rely on the information from $\bm{z}_{t}$ and ignore $\bm{c}_{f}$ during training. Skipping $\bm{z}_{t}$ from ControlNet inputs also helps the Ctrl-Adapter to more effectively handle such sparse frame conditions (see Table 7).
@@
Inverse timestep sampling: robust adaptation to continuous diffusion timestep samplers. While SDv1.5 samples discrete timesteps $t$ uniformly from $\{0,1,...1000\}$, some recent diffusion models [12, 45, 60] sample timesteps from continuous distributions, e.g., SVD [3] samples timesteps from a LogNormal distribution. This gap between discrete and continuous distributions means that we cannot assign the same timestep $t$ to both the video diffusion model and the ControlNet. Therefore, we propose inverse timestep sampling, an algorithm that creates a timestep mapping between the continuous and discrete time distributions (see Algorithm 1 for PyTorch [1] code). The high-level idea of this algorithm is inspired by inverse transform sampling [13]. Given the cumulative distribution functions (CDFs) of the continuous timestep distribution $F_{\text{cont.}}$ and the ControlNet timestep distribution $F_{\text{CNet}}$, we first uniformly sample a value $u$ between $[0,1]$, and then returns the smallest timesteps $t_{\text{cont.}}\in[0,\infty]\subseteq\mathbb{R},t_{\text{CNet}}\in\{0,1,...,1000\}\subseteq\mathbb{N}$, such that $F_{\text{cont.}}(t_{\text{cont.}})\geq u,F_{\text{CNet}}(t_{\text{CNet}})\geq u$. This procedure naturally creates a mapping between two distributions. See Sec. B.2 for details.
Figure 5: Left: Framework for multi-condition video generation by combining multiple ControlNets. $w_{1},w_{2},...,w_{N}$ are the weights allocated to each ControlNet. Right: Three MoE router variants. (a) operates globally, while (b) and (c) operate on the fine-grained patch-level. $C$ and $N$ represent feature dimensions and number of ControlNet experts respectively. $w_{k}^{i,j}$ represents the router weights at position ($i,j$) of the $k^{\text{th}}$ ControlNet 2D feature map. SM stands for Softmax.